# FPGA-Enabled Machine Learning Applications in Earth Observation: A Systematic Review

### v1

@@ 3. Mapping the Research Landscape: Applications, Models, and Hardware
RS Problem Article ML Model FPGA Performance Task Mod. Application Dataset Ref. Year Original Name Core Backbone Family Score [%] MF[MB] C[OP] Classification RGB Landcover/Land use UCM Land Use (Yang and Newsam, 2010) (Zhang et al., 2023b) 2023 A2NN VGG VGG11 Virtex-7 94.76 (OA) - 14.99G NWPU-RESISC45 (Cheng et al., 2017) (Yan et al., 2022) 2022 Improved VGG16 VGG VGG16 Artix-7 88.08 (OA) 14.80 40.96G (Zhang et al., 2020) 2020 Q-IORN VGG VGG16 Virtex-7 88.31 (OA) 121.51 - (Ni et al., 2023) 2023 ResNet-34 ResNet ResNet-34 Virtex-7 92.81 (OA) 21.29 7.33G (Ni et al., 2023) 2023 VGG16 VGG VGG16 Virtex-7 91.90 (OA) 14.70 30.69G UAV RGB (cust.) (Matos-Carvalho et al., 2019) 2019 MLP Shallow NN - Spartan-3A 95.14 (OA) - - Ship id. MASATI (Gallego et al., 2018) (Neris et al., 2022) 2022 AlexNetLite AlexNet AlexNet Kintex US 89 (F1) - 5.71G (Neris et al., 2022) 2022 MobileNetv1Lite Custom CNN MobileNetv1 Kintex US 94 (F1) - 0.05G Kaggle SSI (Ieracitano et al., 2024) 2024 HO-ShipNet Custom CNN - Zynq 7000 93.44 (F1) - - Cloud coverage L8 Biome (Foga et al., 2017) (Pitonak et al., 2022) 2022 CloudSatNet-1 Q2 Custom CNN - Zynq 7000 83.41 (OA) 1.43 - (Pitonak et al., 2022) 2022 CloudSatNet-1 Q4 Custom CNN - Zynq 7000 87.42 (OA) 3.06 - RGB (cust.) (Upadhyay et al., 2024) 2024 ResNet-50 ResNet ResNet-50 Zynq US+ - - - (Upadhyay et al., 2024) 2024 SICNet Custom CNN - Zynq US+ 85.76 (OA) - - Landsat-8 (cust.) (Kim et al., 2024) 2024 TriCloudNet + U-Net U-Net SqueezeNet Zynq 7000 6.84 (CRE) 3.61 - (Vitolo et al., 2024) 2024 CNN Custom CNN - Artix-7 89.64 (F1) - - HSI Air quality mon. UAV RGB (cust.) (Huang et al., 2024) 2024 ETAUS ResNet ResNet-50 Zynq US+ 86.38 (F1) - - Cloud coverage Sentinel-2 (cust.) (Rapuano et al., 2021) 2021 CloudScout Custom CNN - Zynq US+ 92.0 (OA) 13.30 - SAR Military targets id. MSTAR (SDMS, 2010) (Wei et al., 2019) 2019 LeNet-5 f32 LeNet LeNet-5 Kintex-7 98.76 (OA) 6.64 - (Wei et al., 2019) 2019 LeNet-5 i8 LeNet LeNet-5 Kintex-7 97.77 (OA) 1.66 - (Chen et al., 2020b) 2020 LeNet-5 LeNet LeNet-5 Kintex-7 98.18 (OA) - - (Zhang et al., 2022) 2022 GNN GNN - Zynq US+ 99.09 (OA) 0.96 2.13M (Zhang et al., 2023a) 2023 GNN GNN - Alveo - - - Ship id. ALOS-2 (cust.) (Hashimoto et al., 2019) 2019 CNN Custom CNN - Zynq 7000 100 (OA) - - (Myojin et al., 2020) 2020 BNN Custom CNN - Zynq 7000 - - - Pixel classification RGB Deforestation det. UAV RGB (cust.) (Torres et al., 2020) 2020 Weightless Neural Systems Shallow NN - Zynq 7000 90.0 (OA) - - UAV landing site id. UAV RGB (cust.) (Fraczek et al., 2018) 2018 Decision Tree Trad. ML - Zynq 7000 92.1 (OA) - - Landcover/Land use ALSAT-2A (cust.) (Yahiaoui et al., 2017) 2017 BRAM_DSP Trad. ML - Virtex-6 94.8 (OA) - - (Yahiaoui et al., 2017) 2017 LUT_MUL Trad. ML - Virtex-6 94.8 (OA) - - HSI Landcover/Land use University of Pavia (UPV/EHU, 2021) (Martins et al., 2024) 2024 SVM Trad. ML - Zynq 7000 82.48 (AA) - - (He et al., 2023) 2023 2D CNN Custom CNN - Zynq US+ 98.24 (OA) 1.20 0.684M (He et al., 2023) 2023 3D CNN Custom CNN - Zynq US+ 94.09 (OA) 0.12 4.231M (He et al., 2023) 2023 HybridSN Custom CNN - Zynq US+ 100 (OA) 20.50 101.642M (Shibi and Gayathri, 2021) 2021 LPDBL Trad. ML - Virtex-7 - - - (Chellaswamy et al., 2024) 2024 SAM-GNN GNN - Virtex-7 95.05 (OA) - - AVIRIS-NG (Gyaneshwar and Nidamanuri, 2022) 2022 CAG-SC2S Trad. ML - Artix-7 98.30 (OA) - - (Gyaneshwar and Nidamanuri, 2022) 2022 CAL-SC2S Trad. ML - Artix-7 99.69 (OA) - - (Gyaneshwar and Nidamanuri, 2022) 2022 SVM Trad. ML - Artix-7 20.71 (OA) - - Anomaly det. ABU (Kang, 2017) (Boyle et al., 2023) 2023 Deep Belief Network Shallow NN - Zynq US+ - - - Segmentation RGB Urban areas Potsdam (ISPRS, 2022) (Sabogal and George, 2021) 2021 ENet Custom CNN - Zynq US+ 63.3 (mIoU) 0.36 4.06G (Sabogal and George, 2021) 2021 ESPNet Custom CNN - Zynq US+ 55.5 (mIoU) 0.33 3.71G (Sabogal and George, 2021) 2021 FPN Custom CNN - Zynq US+ 65.3 (mIoU) 5.84 17.30G (Sabogal and George, 2021) 2021 U-Net U-Net - Zynq US+ 61.4 (mIoU) 7.40 96.68G General Landsat-8-OLI (cust.) (Ratnakumar and Nanda, 2021) 2021 RDBC Trad. ML - Virtex-6 37.45 (SSIM) - O(n) Cloud coverage 38-Cloud (Mohajerani, 2024) (Bahl et al., 2019) 2019 C-FCN++ Custom CNN - Cyclone V 79.30 (mIoU) 0.047 4.7K Object detection RGB Diverse det. DOTAv1.0 (Xia et al., 2018) (Zhang et al., 2021) 2021 Improved YOLOv2 YOLO Darknet19 Zynq 7000 67.32 (mAP) - 379G (Yan et al., 2022) 2022 Improved YOLOv2 YOLO Darknet19 Artix-7 67.30 (mAP) 49.40 379.55G (Ni et al., 2023) 2023 Improved YOLOv2 YOLO Darknet19 Virtex-7 67.30 (mAP) 49.40 379.55G (Yang et al., 2023) 2023 Ghost-YOLOS YOLO GhostNet Zynq 7000 62.58 (mAP) 12.60 8.41G DIOR (Li et al., 2020b) (Zhao et al., 2023) 2023 YOLOv4-MobileNetv3 YOLO MobileNetv3 Zynq US+ 82.61 (mAP) 5.69 - (Li et al., 2023) 2023 RFA-YOLO YOLO MobileNeXt Zynq US+ 64.85 (mAP) 24.75 - NWPU VHR-10 (Cheng et al., 2014) (Li et al., 2019) 2019 CBFF-SSD SSD MobileNetv1 Zynq 7000 91.42 (mAP) - - DAC 2018 (Wang and Qiu, 2019) 2019 CNN Custom CNN - Zynq 7000 21 (mIoU) - - UAV RGB (cust.) (Li et al., 2020a) 2020 AP2D-Net Custom CNN - Zynq US+ 55.0 (mIoU) - - Flying-object det. PennSyn2Real (Nguyen et al., 2021) (Suh et al., 2021) 2021 SSD 0.25x SSD VGG16 Zynq US+ 76.2 (mAP) - 4.04G (Suh et al., 2021) 2021 SSD 0.5x SSD VGG16 Zynq US+ 83.9 (mAP) - 15.65G (Suh et al., 2021) 2021 SSD 1.0x SSD VGG16 Zynq US+ 88.4 (mAP) - 61.60G UAV RGB (cust.) (Nguyen et al., 2024) 2024 YOLOv4 YOLO Darknet53 Zynq US+ 83.7 (mAP) 245 105.9G (Nguyen et al., 2024) 2024 YOLOv4-tiny 3L YOLO Darknet53-tiny Zynq US+ 72.24 (mAP) 24.10 15.32G Aircraft det. Airbus (Garg et al., 2024) 2024 ResNet-18+YOLOv2 YOLO ResNet-18 Zynq US+ 91 (AP) - - GES RGB (sim.) (Wu et al., 2021) 2021 Improved YOLOv3 YOLO DarkNet53 Zynq US+ 92 (OA) 8.50 - UAV obstacles det. UAV RGB+MMW (cust.) (Wang et al., 2024) 2024 Lightweight CNN ResNet ResNet-18 Zynq 7000 70.1 (mAP) - 4.6G Railway defect det. FastenerDataset (cust.) (Yu et al., 2024) 2024 Improved YOLOv4-tiny YOLO Darknet53-tiny Zynq US+ 95.1 (mAP) - - SAR Ship det. SSDD (Li et al., 2017) (Yang et al., 2022) 2022 CNN2@0.7 YOLO MobileNetv1 Virtex-7 94 (AP) 0.09 0.45G (Yang et al., 2022) 2022 CNN4@0.7 YOLO MobileNetv2 Virtex-7 93.3 (AP) 0.11 0.56G (Yang et al., 2022) 2022 CNN6@1.6 YOLO SqueezeNet Virtex-7 92.8 (AP) 0.09 0.6G Regr. SAR Oil spills mon. SAR (sim.) (Hammoud et al., 2022) 2022 MLP Shallow NN - Zynq 7000 - - - Task: Regr. Regression; Dataset: GES Google Earth Studio, MMW MilliMeter-Wave radar; Model: RDBC Roller Dung Bettle Clustering, WNS Weightless Neural System
@@ 3.1. Remote Sensing/Machine Learning Taxonomy
To help readers explore studies addressing similar problem settings, we group the collected studies by how RS applications are formulated as ML tasks. Table 2 presents this taxonomy, clustering all reported experiments across the selected studies. When a study conducts several relevant experiments, we extract each corresponding model and metric. For example, Suh et al. (2021) present a methodological contribution and evaluate several sizes of the same architecture, i.e., SSD 0.25x, SSD 0.5x, and SSD 1.0x. We argue that reporting each experiment separately, rather than considering a single experiment per study, portrays a more complete picture of the conducted work. In total, the taxonomy includes $66$ experiments drawn from $46$ studies.
Each row in Table 2 corresponds to an experiment. The first group of columns, dedicated to the RS Problem, introduces the taxonomy: experiments are hierarchically grouped by ML ’Task’, i.e., the family of the ML problem, data modality (’Mod.’), i.e., the type of RS data used as input, ’Application’ and ’Dataset’. As a result, experiments from the same study may not be juxtaposed. ’Application’ refers to the downstream task tackled by the study; applications are closely tied with the ’Dataset’ [^5] Datasets carrying the (cust.) tag are unpublished datasets, meaning that the authors built their dataset but did not name or publish it. and are detailed later in Fig. 1. The ’Article’ group provides context: ’Ref.’ lists the source article and ’Year’ indicates its publication date.
The rest of the table is dedicated to the ML component of the experiment. ’ML Model’ characterizes the model’s architecture through three fields: ’Original Name’, the exact appellation used in the study; Core, a high-level architectural descriptor (e.g., CNN, Shallow NN such as a Multi-Layer Perceptron (MLP), or Traditional ML like an SVM), further detailed in Fig. 2; and, for Deep Neural Networks (DNNs), their ’Backbone’. Next, the ’FPGA’ column indicates the family of the FPGA device used for implementation, more details in Fig. 3. The last group of columns ’Performance’ reports a few metrics of the experiment: ’Score’ corresponds to the accuracy in [%] of the model on its dataset and ’MF[MB] is the Memory Footprint of the model in MegaBytes. Finally, ’C[OP]’ stands for the computational complexity of the model in numbers of operations [^6] More details about the distinction between [GOP/s], [GOP], or [GOPs] can be found in Section 6.3: Recommendations.. These performance metrics are the most closely related to the ML model; further metrics are reported during the second literature taxonomy in Table 3.
@@
From urban planning to disaster assessment and target monitoring, Remote Sensing images support a wide and growing range of Earth Observation applications (Zhao et al., 2022).
[…]
[Bubble chart grouping RS applications by thematics] A bubble chart showing applications of Remote Sensing data, colored by corresponding formulation as an ML task. The main categories are Environmental Monitoring, Landscape Analysis, Target Surveillance, and Navigation, represented by large transparent circles containing applications denoted by smaller circular patches. Environmental monitoring includes applications like deforestation detection, air quality monitoring, oil spills feature extraction, and cloud coverage, mostly using classification and regression techniques. Landscape Analysis encompasses anomaly detection, urban areas, unsupervised learning, and landcover/land use, employing pixel classification and segmentation. Target Surveillance addresses railway defect detection, aircraft detection, military target identification, ship identification, flying-object detection, and miscellaneous targets mainly through object detection and classification techniques. Navigation is the smallest circle, it involves safe UAV landing site identification and UAV obstacle detection. The chart also features a legend indicating which remote sensing data sources (RGB, HSI, SAR) are used in each application.
Figure 1. Addressed RS applications grouped by thematic and colored as their corresponding ML task. The size of the circular patches represents the number of experiments, while the hatching pattern represents the prevailing data modality.
@@
Fig. 1 separates the surveyed RS applications into four distinct thematics. Target Surveillance [^7] Target detection is a more frequent appellation for RS applications. However, we will prefer the term ”surveillance” to avoid any confusion with the ML task ”object detection,” sometimes also called ”target detection.” gathers all applications focused on recognizing a specific family of objects. The output of such a task can be the location—bounding box—of the targets, in such a case we denote it detection (det.), or the binary presence of the target in the observed image, denoted identification (id.). In this thematic, ”Miscellaneous targets” refers to heterogeneous object families, such as ships and cars, but also tennis courts or roundabouts in the DOTAv1.0 (Xia et al., 2018) or DIOR (Li et al., 2020b) datasets. Landscape analysis applications investigate the type of terrain present in the image. In this survey, ”Urban areas” refers to segmenting city tiles, e.g., with the POTSDAM dataset (ISPRS, 2022). ”Landcover/Land use” is synonym to the classification of the type of terrain, such as desert, stadium, sea ice, etc., for example, in NWPU-RESISC45 (Cheng et al., 2017). Finally, Environment monitoring includes landscape analysis applications specifically focused on environmental phenomena, such as cloud or deforestation detection. While the grouping of Fig 1 offers a clear structure for the surveyed studies, it does not capture the full range of RS applications. In particular, the records miss prevalent applications such as natural disaster monitoring or studies of the climate crisis.
@@
The hash pattern in Fig. 1 represents the modality of the images used for each application. In RS, the modality refers to the type of sensor that acquires the data. The most common modality, as in this survey, is optical data, with the three traditional Red-Green-Blue (RGB) channels. In RS, channels are commonly referred to as bands. Indeed, sensors capture data around specific wavelengths, each wavelength’s spectrum resulting in a different band. Hyperspectral Imaging (HSI) emerges as the most extreme case, sometimes with over 200 bands, providing a very fine-grained spectral resolution that approximates a contiguous spectrum (Lopez et al., 2013). Lastly, Synthetic Aperture Radar (SAR) is an active sensor, in opposition to passive sensors that collect light from the observed scene. SAR emits waves that scatter upon hitting the observed scene, then captures the reflected components that bounce back towards the sensor. Such a system allows to collect data during the night and through clouds, making SAR a versatile modality.
@@
In Fig. 1, each application circle is colored depending on the ML task it is primarily formulated as. An ML task is the type of prediction made by the model. The most traditional tasks are Classification and Regression. Classification assigns a category to each input image, while Regression predicts a numerical value from their features. Beyond these, a Clustering task groups data instances into clusters based on shared characteristics. When applied to images, Clustering yields a segmentation map, leading us to prefer the term Segmentation. Pixel Classification also produces a segmentation map by individually processing and categorizing each pixel. However, Pixel Classification typically uses pixel-level features (e.g., spectral values, texture) as input, whereas Segmentation also considers contextual information and spatial relationships between pixels. Finally, Object Detection not only assigns categories but also provides bounding boxes indicating the objects’ location, enabling the identification and categorization of multiple objects within a single image. While some applications are closely associated with a specific ML task, others are more versatile and can be formulated in diverse ways. Fig. 1 includes two examples: ”Landcover/Land use,” formulated five times as a Pixel Classification task and four times as a Classification task, and ”Cloud Coverage Classification,” expressed once as a Segmentation task. Ultimately, how RS applications are formulated as ML tasks reflects the decisions and priorities of the individual studies.

### v2

@@ 3. Mapping the Research Landscape: Applications, Models, and Hardware
ML Remote Sensing Article ML Model FPGA Performance Task Mod. Application Dataset Ref. Year Original Name Core Backbone Family Score [%] MF[MB] C[OP] Classification RGB Cloud coverage class. L8 Biome (Foga et al., 2017) (Pitonak et al., 2022) 2022 CloudSatNet-1 Q2 Custom CNN - Zynq 7000 83.41 (OA) 1.43 - (Pitonak et al., 2022) 2022 CloudSatNet-1 Q4 Custom CNN - Zynq 7000 87.42 (OA) 3.06 - Sentinel-2 (cust.) (Rapuano et al., 2021) 2021 CloudScout Custom CNN - Zynq US+ 92.0 (OA) 13.30 - RGB (cust.) (Upadhyay et al., 2024) 2024 ResNet-50 ResNet ResNet-50 Zynq US+ - - - (Upadhyay et al., 2024) 2024 SICNet Custom CNN - Zynq US+ 85.76 (OA) - - Landsat-8 (cust.) (Vitolo et al., 2024) 2024 CNN Custom CNN - Artix-7 89.64 (F1) - - Land Use Land Cover UCM Land Use (Yang and Newsam, 2010) (Zhang et al., 2023) 2023 A2NN VGG VGG11 Virtex-7 94.76 (OA) - 14.99G NWPU-RESISC45 (Cheng et al., 2017) (Zhang et al., 2020) 2020 Q-IORN VGG VGG16 Virtex-7 88.31 (OA) 121.51 - (Yan et al., 2022) 2022 Improved VGG16 VGG VGG16 Artix-7 88.08 (OA) 14.80 40.96G (Ni et al., 2023) 2023 ResNet-34 ResNet ResNet-34 Virtex-7 92.81 (OA) 21.29 7.33G (Ni et al., 2023) 2023 VGG16 VGG VGG16 Virtex-7 91.90 (OA) 14.70 30.69G UAV RGB (cust.) (Matos-Carvalho et al., 2019) 2019 MLP Shallow NN - Spartan-3A 95.14 (OA) - - Ship id. MASATI (Gallego et al., 2018) (Neris et al., 2022) 2022 AlexNetLite AlexNet AlexNet Kintex US 89 (F1) - 5.71G (Neris et al., 2022) 2022 MobileNetv1Lite Custom CNN MobileNetv1 Kintex US 94 (F1) - 0.05G Kaggle SSI (Ieracitano et al., 2024) 2024 HO-ShipNet Custom CNN - Zynq 7000 93.44 (F1) - - SAR Military targets id. MSTAR (SDMS, 2010) (Wei et al., 2019) 2019 LeNet-5 f32 LeNet LeNet-5 Kintex-7 98.76 (OA) 6.64 - (Wei et al., 2019) 2019 LeNet-5 i8 LeNet LeNet-5 Kintex-7 97.77 (OA) 1.66 - (Chen et al., 2020b) 2020 LeNet-5 LeNet LeNet-5 Kintex-7 98.18 (OA) - - (Zhang et al., 2022) 2022 GNN GNN - Zynq US+ 99.09 (OA) 0.96 2.13M (Zhang et al., 2023) 2023 GNN GNN - Alveo - - - (Wickramasinghe et al., 2024) 2024 VTR ViT - Alveo 95.96 (OA) - 18.28G Ship id. ALOS-2 (cust.) (Hashimoto et al., 2019) 2019 CNN Custom CNN - Zynq 7000 100 (OA) - - (Myojin et al., 2020) 2020 BNN Custom CNN - Zynq 7000 - - - HSI Air quality class. UAV RGB (cust.) (Huang et al., 2024) 2024 ETAUS ResNet ResNet-50 Zynq US+ 86.38 (F1) - - Segmentation - Pixel RGB Deforestation det. UAV RGB (cust.) (Torres et al., 2020) 2020 Weightless Neural Systems Shallow NN - Zynq 7000 90.0 (OA) - - UAV landing site id. UAV RGB (cust.) (Fraczek et al., 2018) 2018 Decision Tree Trad. ML - Zynq 7000 92.1 (OA) - - RGB Land Use Land Cover ALSAT-2A (cust.) (Yahiaoui et al., 2017) 2017 BRAM_DSP Trad. ML - Virtex-6 94.8 (OA) - - (Yahiaoui et al., 2017) 2017 LUT_MUL Trad. ML - Virtex-6 94.8 (OA) - - HSI Land Use Land Cover University of Pavia (Shibi and Gayathri, 2021) 2021 LPDBL Trad. ML - Virtex-7 - - - (He et al., 2023) 2023 2D CNN Custom CNN - Zynq US+ 98.24 (OA) 1.20 0.684M (He et al., 2023) 2023 3D CNN Custom CNN - Zynq US+ 94.09 (OA) 0.12 4.231M (He et al., 2023) 2023 HybridSN Custom CNN - Zynq US+ 100 (OA) 20.50 101.642M (Martins et al., 2024) 2024 SVM Trad. ML - Zynq 7000 82.48 (AA) - - (Chellaswamy et al., 2024) 2024 SAM-GNN GNN - Virtex-7 95.05 (OA) - - AVIRIS-NG (Gyaneshwar and Nidamanuri, 2022) 2022 CAG-SC2S Trad. ML - Artix-7 98.30 (OA) - - (Gyaneshwar and Nidamanuri, 2022) 2022 CAL-SC2S Trad. ML - Artix-7 99.69 (OA) - - (Gyaneshwar and Nidamanuri, 2022) 2022 SVM Trad. ML - Artix-7 20.71 (OA) - - Land cover change det. ABU (Kang, 2017) (Boyle et al., 2023) 2023 Deep Belief Network Shallow NN - Zynq US+ - - - Segmentation - Patch RGB Urban areas Potsdam (ISPRS, 2022) (Sabogal and George, 2021) 2021 ENet Custom CNN - Zynq US+ 63.3 (mIoU) 0.36 4.06G (Sabogal and George, 2021) 2021 ESPNet Custom CNN - Zynq US+ 55.5 (mIoU) 0.33 3.71G (Sabogal and George, 2021) 2021 FPN Custom CNN - Zynq US+ 65.3 (mIoU) 5.84 17.30G (Sabogal and George, 2021) 2021 U-Net U-Net - Zynq US+ 61.4 (mIoU) 7.40 96.68G Cloud extr. 38-Cloud (Mohajerani and Saeedi, 2019) (Bahl et al., 2019) 2019 C-FCN++ Custom CNN - Cyclone V 79.30 (mIoU) 0.047 4.7K Landsat-8 (cust.) (Kim et al., 2024) 2024 TriCloudNet + U-Net U-Net SqueezeNet Zynq 7000 6.84 (CRE) 3.61 - Land Use Land Cover Landsat-8-OLI (cust.) (Ratnakumar and Nanda, 2021) 2021 RDBC Trad. ML - Virtex-6 37.45 (SSIM) - O(n) Object detection RGB Diverse det. DOTAv1.0 (Xia et al., 2018) (Zhang et al., 2021) 2021 Improved YOLOv2 YOLO DarkNet19 Zynq 7000 67.32 (mAP) - 379G (Yan et al., 2022) 2022 Improved YOLOv2 YOLO DarkNet19 Artix-7 67.30 (mAP) 49.40 379.55G (Yang et al., 2023) 2023 Ghost-YOLOS YOLO GhostNet Zynq 7000 62.58 (mAP) 12.60 8.41G (Ni et al., 2023) 2023 Improved YOLOv2 YOLO DarkNet19 Virtex-7 67.30 (mAP) 49.40 379.55G DIOR (Li et al., 2020a) (Zhao et al., 2023) 2023 YOLOv4-MobileNetv3 YOLO MobileNetv3 Zynq US+ 82.61 (mAP) 5.69 - (Li et al., 2023a) 2023 RFA-YOLO YOLO MobileNeXt Zynq US+ 64.85 (mAP) 24.75 - NWPU VHR-10 (Cheng et al., 2014) (Li et al., 2019b) 2019 CBFF-SSD SSD MobileNetv1 Zynq 7000 91.42 (mAP) - - DAC 2018 (Wang and Qiu, 2019) 2019 CNN Custom CNN - Zynq 7000 21 (mIoU) - - UAV RGB (cust.) (Li et al., 2020b) 2020 AP2D-Net Custom CNN - Zynq US+ 55.0 (mIoU) - - Flying-object det. PennSyn2Real (Nguyen et al., 2021) (Suh et al., 2021) 2021 SSD 0.25x SSD VGG16 Zynq US+ 76.2 (mAP) - 4.04G (Suh et al., 2021) 2021 SSD 0.5x SSD VGG16 Zynq US+ 83.9 (mAP) - 15.65G (Suh et al., 2021) 2021 SSD 1.0x SSD VGG16 Zynq US+ 88.4 (mAP) - 61.60G UAV RGB (cust.) (Nguyen et al., 2024) 2024 YOLOv4 YOLO DarkNet53 Zynq US+ 83.7 (mAP) 245 105.9G (Nguyen et al., 2024) 2024 YOLOv4-tiny 3L YOLO DarkNet53-tiny Zynq US+ 72.24 (mAP) 24.10 15.32G Aircraft det. Airbus (Garg et al., 2024) 2024 ResNet-18+YOLOv2 YOLO ResNet-18 Zynq US+ 91 (AP) - - GES RGB (sim.) (Wu et al., 2021) 2021 Improved YOLOv3 YOLO DarkNet53 Zynq US+ 92 (OA) 8.50 - Railway defect det. FastenerDataset (cust.) (Yu et al., 2024) 2024 Improved YOLOv4-tiny YOLO DarkNet53-tiny Zynq US+ 95.1 (mAP) - - RGB UAV obstacles det. UAV RGB+MMW (cust.) (Wang et al., 2024) 2024 Lightweight CNN ResNet ResNet-18 Zynq 7000 70.1 (mAP) - 4.6G SAR Ship det. SSDD (Li et al., 2017) (Yang et al., 2022) 2022 CNN2@0.7 YOLO MobileNetv1 Virtex-7 94 (AP) 0.09 0.45G (Yang et al., 2022) 2022 CNN4@0.7 YOLO MobileNetv2 Virtex-7 93.3 (AP) 0.11 0.56G (Yang et al., 2022) 2022 CNN6@1.6 YOLO SqueezeNet Virtex-7 92.8 (AP) 0.09 0.6G Regr. SAR Oil spills feature extr. SAR (sim.) (Hammoud et al., 2022) 2022 MLP Shallow NN - Zynq 7000 - - - (Hammoud et al., 2024) 2024 TinyML Shallow NN - Zynq 7000 99.00 (OA) 0.000037 - Task: Regr. Regression; Dataset: GES Google Earth Studio, MMW MilliMeter-Wave radar; Model: RDBC Roller Dung Bettle Clustering, WNS Weightless Neural System Modality: $\dagger$ = RGB + infrared; $\ddagger$ = RGB + MMW
@@ 3.1. Remote Sensing/Machine Learning Taxonomy
To help readers explore studies addressing similar problem settings, we present in Table 2 a taxonomy grouping the collected studies based on how RS applications are formulated as ML tasks. When a study conducts several relevant experiments, we extract each corresponding model and metric. For example, Suh et al. (2021) present a methodological contribution and evaluate several sizes of the same architecture, i.e., SSD 0.25x, SSD 0.5x, and SSD 1.0x. We argue that reporting each experiment separately, rather than considering a single experiment per study, portrays a more complete picture of the conducted work. In total, the taxonomy includes $68$ experiments drawn from $48$ studies.
Each row in Table 2 corresponds to an experiment. In the first column, we classify the ML problem into different Computer Vision (CV) tasks. To facilitate comparison with other ML works, we describe the ML ’Task’ using CV terminology over RS nomenclature [^5] For example, we use the term Classification for per-image class assignment while it implies a per-pixel class assignment in RS.. In addition, we distinguish between Segmentation models that process a pixel at a time (’Segmentation - Pixel’) and those that process the patch as a whole (’Segmentation - Patch’), as the throughput requirements of the latter are significantly higher. Compared to the first column, the rest of the taxonomy uses RS terminology, as it focuses on the goals of the experiments, i.e., the RS-relevant information they generate. In the following group of columns, dedicated to the RS Problem, experiments are hierarchically grouped by data modality (’Mod.’), i.e., the type of RS data used as input, ’Application’, and ’Dataset’. As a result, experiments from the same study may not be juxtaposed. ’Application’ refers to the downstream task tackled by the study; applications are closely tied with the ’Dataset’ [^6] Datasets carrying the (cust.) tag are unpublished datasets, meaning that the authors built their dataset but did not name or publish it. and are detailed later in Fig. 2. The ’Article’ group provides the experiment’s context: ’Ref.’ lists the source article and ’Year’ indicates its publication date.
The rest of the table is dedicated to the ML component of the experiment. ’ML Model’ characterizes the model’s architecture through three fields: ’Original Name’, the exact appellation used in the study; Core, a high-level architectural descriptor (e.g., CNN, Shallow NN such as a Multi-Layer Perceptron (MLP), or Traditional ML like an SVM), further detailed in Fig. 3; and, for Deep Neural Networks (DNNs), their ’Backbone’, i.e., the core of the network responsible for extracting features. Next, the ’FPGA’ column indicates the family of the FPGA device used for implementation, more details in Fig. 4. The last group of columns ’Performance’ gathers a few metrics of the experiment as reported in the original study: ’Score’ corresponds to the accuracy in [%] of the model on its dataset and ’MF[MB] is the Memory Footprint of the model in MegaBytes. Finally, ’C[OP]’ stands for the computational complexity of the model in numbers of operations. These performance metrics are the most closely related to the ML model; further metrics are reported during the second literature taxonomy in Table 4.
@@
From disaster assessment to target monitoring, RS images support a wide and growing range of EO applications (Zhao et al., 2022).
Refer to caption Bubble chart grouping RS applications by thematics A bubble chart showing applications of Remote Sensing data, colored by corresponding formulation as an ML task. The main categories are Land & Environment, Weather & Atmosphere, Target Surveillance, and UAV Navigation, represented by large transparent circles containing applications denoted by smaller circular patches. Land & Environment includes several segmentation tasks for urban Areas or Land Use Land Cover, but also railway defects detection or oil spills feature extraction. Weather & Atmosphere applications classify Cloud coverage or air quality, or extract clouds from images. Target Surveillance addresses aircraft detection, military target identification, ship identification or detection, flying-object detection, and miscellaneous targets mainly through object detection and classification techniques. UAV Navigation is the smallest circle; it involves segmenting safe UAV landing sites and detecting UAV obstacles. The chart also features a legend indicating which remote sensing data modalities (RGB, HSI, SAR) are used in each application.
Figure 2. Addressed RS applications grouped by thematic and colored as their corresponding ML task. The circles’ size represents the number of experiments, while the hatching pattern represents the prevailing data modality.
@@
Fig. 2 separates the surveyed RS applications into four distinct thematics. Target Surveillance [^7] Target detection is a more frequent appellation for RS applications. However, we will prefer the term ”surveillance” to avoid any confusion with the ML task ”object detection,” sometimes also called ”target detection.” gathers all applications focused on recognizing a specific family of objects. The output of such a task can be the location—bounding box—of the targets, in such a case we denote it detection , or the binary presence of the target in the observed image, denoted identification . In this thematic, ”Miscellaneous targets” refers to heterogeneous object families, such as ships and cars, but also tennis courts or roundabouts in the DOTAv1.0 (Xia et al., 2018) or DIOR (Li et al., 2020a) datasets. ”Flying-object detection” and ”Aircraft detection” differentiate objects in the air (planes, drones, birds, etc.) and standing aircraft on the ground. Land & Environment applications monitor the type of terrain present in the image, either on a general level, such as Land Use Land Cover (LULC), or for specific phenomena, such as oil spills. In particular, ”Land cover change detection” performs semantic segmentation of the land in an unsupervised fashion to identify pixels with abnormal values. In Weather & Atmosphere, we distinguish experiments that extract cloud masks from images and experiments that classify complete scenes based on cloud coverage thresholds. Finally, UAV Navigation applications rapidly use the acquired images to support the UAV’s itinerary. While the grouping of Fig 2 offers a clear structure for the surveyed studies, it does not capture the full range of RS applications. In particular, the records miss prevalent applications such as natural disaster monitoring or studies of the climate crisis.
@@
The hash pattern in Fig. 2 represents the modality of the images used for each application. In RS, the modality refers to the type of sensor used to acquire the data. The most common modality, as in this survey, is optical data [^8] Optical data is commonly named Multispectral in RS. However, as $\sim\!95\%$ of the surveyed studies only use the RGB channels of the electromagnetic spectrum, we use the term optical or RGB for accuracy., with the three traditional Red-Green-Blue (RGB) channels. In RS, channels are commonly referred to as bands. Indeed, sensors capture data around specific wavelengths, each wavelength’s spectrum resulting in a different band. Hyperspectral Imaging (HSI) emerges as the most extreme case, sometimes with over 200 bands, providing a very fine-grained spectral resolution that approximates a contiguous spectrum (Lopez et al., 2013). Lastly, Synthetic Aperture Radar (SAR) is an active sensor, in opposition to passive sensors that collect light from the observed scene. SAR emits waves that scatter upon hitting the observed scene, then captures the reflected components that bounce back towards the sensor. Such a system enables data collection during nighttime and through clouds, making SAR a versatile modality.
@@
In Fig. 2, each application circle is colored depending on the ML task it is primarily formulated as. An ML task is the type of prediction made by the model. The most traditional tasks are Classification and Regression. Also referred to as Scene Classification and Parameter Retrieval in RS terminology, the former assigns a category to each input image, while the latter predicts a numerical value from its features. Beyond these, a Clustering task groups data instances into clusters based on shared characteristics. When applied to images, Clustering yields a segmentation map, leading us to prefer the term Segmentation. Semantic segmentation maps can be generated by processing the entire input image as a whole or by processing pixels one by one repetitively across the image. Finally, Object Detection not only assigns categories but also provides bounding boxes indicating the location of objects, enabling the identification and categorization of multiple objects within a single image. While some applications are closely tied to a specific ML task, others are more versatile and can be formulated in various ways. Fig. 2 includes one example: ”LULC”, was formulated once as Segmentation - Patch, six times as Classification, and 12 times as a Segmentation - Pixel task. Ultimately, how RS applications are formulated as ML tasks reflects the decisions and priorities of each study.